please help me with this...
12p - 6.5p + 7 = -15

will you show me how you did it?

Answer by rahman(247)

12p - 6.5p + 7 = -15 ------->multiply by 2
24p - 13p + 14 = -30
(24 - 13)p = -30 - 14
11p = - 44 -----------> Multiply by 11
p = -4
OK